OpenStudy (anonymous):

medal if helped Choose one of the factors of 1331x3 - 8y3. 4 11x - 2y 121x2 + 44xy + 4y2 121x2 - 22xy + 4y2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

a little help plz

OpenStudy (amistre64):

theres a formula that defines a difference of cubes

OpenStudy (anonymous):

i think its...

OpenStudy (anonymous):

a3 + b3 = (a + b)(a2 – ab + b2) a3 – b3 = (a – b)(a2 + ab + b2)

OpenStudy (amistre64):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

a is 1331 right and b is 8

OpenStudy (amistre64):

a is 11x and b is 2y

OpenStudy (anonymous):

wah

OpenStudy (anonymous):

oh cuz u cube rooted them i see

OpenStudy (amistre64):

(11x)^3 = 1331x^3 (2y)^3 = 8y^3

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so it would b 11x-2y?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

ok i get thx

OpenStudy (anonymous):

medal 4 u

OpenStudy (amistre64):

the other possible one is C, but 8*11 = 88 so thats just not going to work

OpenStudy (anonymous):

yeah i thought it was c at first thats y i asked

OpenStudy (amistre64):

:) good luck